OpenStudy (anonymous):

find dy/dx of y = (2x^7-x^2) (x-1/x+1)

OpenStudy (helder_edwin):

u have \[ \large (2x^7-x^2)\cdot\left(\frac{x-1}{x+1}\right) \] do u know which rules to use?

OpenStudy (anonymous):

umm do you use the Quotient rule on the (x-1/x+1) right?

OpenStudy (anonymous):

then product rule?

OpenStudy (helder_edwin):

actually, first the product rule and then the quotient rule

OpenStudy (helder_edwin):

u have \[\large (AB)'=A'B+AB' \] but B is quotient so \[\large B=\frac{C}{D} \] then \[\large B'=\frac{C'D-CD'}{D^2} \] got it?

OpenStudy (anonymous):

kinda..

OpenStudy (helder_edwin):

try it!

OpenStudy (helder_edwin):

post your solution, so we can help you better

OpenStudy (anonymous):

hmm can you explain the above thing better ^

OpenStudy (helder_edwin):

your function has two "main" operations, one multiplication and one division. right? these tell which of the rules of differentiation to use. now since multiplication comes first (is to the left of the division) u apply this rule first. ok?

OpenStudy (helder_edwin):

u have \[\large y=(2x^7-x^2)\cdot\left(\frac{x-1}{x+1}\right) \] so \[\large y'=(2x^7-x^2)'\left(\frac{x-1}{x+1}\right)+(2x^7-x^2) \left(\frac{x-1}{x+1}\right)' \] ok?

OpenStudy (anonymous):

ohhh okay i see

OpenStudy (helder_edwin):

try it! u have to try.

OpenStudy (anonymous):

okay so do i do the Quotient rule now?

OpenStudy (helder_edwin):

u have two derivatives, for the second yes, u have to do the quotient rule